OpenStudy (anonymous):

I have been having a lot of trouble with this one. Please help! 6/(5s^2-5s) - 5/(5s-5) subtract. simplify if possible.

OpenStudy (saifoo.khan):

\[\huge \frac{6}{5s^2 -5s}-\frac{5}{5s-5}\]

OpenStudy (saifoo.khan):

Please Confirm! \[\huge \frac{6-5s}{2s(s-1)}\]

OpenStudy (zarkon):

why the 2 in the denominator?

OpenStudy (anonymous):

it must be a typo

OpenStudy (saifoo.khan):

ummm, no not typo.

OpenStudy (anonymous):

because it is squared. it is not a typo

OpenStudy (zarkon):

\[\frac{6-5s}{5s(s-1)}\]